# Graph Learning

### v1

@@ 7.1.1 Differential Privacy Based Methods
xiang2024preserving present a solution that focuses on node-level privacy through a two-component protocol: 1) Heter-Poisson, a specialized node sampling strategy that generates sub-graphs with desired properties, and 2) a randomization routine utilizing symmetric multivariate Laplace noise, offering a robust privacy guarantee. Additionally, the authors critiqued a prior approach [sajadmanesh2023gap] that utilizes differentially private node embeddings, highlighting fundamental privacy flaws and proving the impossibility of achieving both strong privacy and acceptable utility with such embeddings, thus underscoring intrinsic utility barriers in enforcing differential privacy.
@@ 7.2.1 Group Fairness
GNNs can produce biased predictions based on protected sensitive attributes, such as skin color and gender, due to historical biases in the training data. This discrimination can be exacerbated by the graph structures and the message-passing mechanisms inherent in GNNs, particularly in high-stakes applications like crime rate prediction. While extensive research has focused on fair classification with independently and identically distributed (i.i.d.) data, approaches for non-i.i.d. scenarios remain limited, especially since sensitive attributes are often scarce in real-world graphs. To tackle this issue, FairGNN [dai2023learning] has been proposed to mitigate bias while maintaining node classification accuracy by leveraging limited sensitive information, and it employs local differential privacy to protect sensitive attributes. An extension, NT-FairGNN, further addresses the challenges of achieving fairness and privacy simultaneously, with theoretical analysis and empirical results demonstrating their effectiveness.
@@ 8.1 Graph Foundation Models
Graph Foundation Models (GFMs) is a large-scale pre-trained model specifically designed for processing graph-structured data. These models are generally pre-trained on vast amounts of data across various domains [xia2022asurvey, lengthiu2023towardsgraph]. Besides, GFMs can adapt to varying demands with minimal fine-tuning or even in few-shot/zero-shot scenarios. There exist many open issues to be addressed, and we discuss them from three distinct perspectives: pre-training, architecture design, and adaptation strategies.
@@ 8.1.2 Architecture Design
  • Message-Passing Models: Message-passing models are the most common designs in graph models. These architectures capture local graph structures by iteratively passing and aggregating information between nodes. Classic models in this category include GCN [KipfW17gcn], GAT [velickovic2018graph], and GraphSAGE [hamiltonYL17graphsage].
  • Graph Transformer: Inspired by the success of transformers in natural language processing, graph transformers have been extended to graph-structured data. These models apply global attention mechanisms to capture long-range dependencies in graphs, making them well-suited for handling complex relations.
  • Generative Models: Graph autoencoders and generative models focus on learning latent representations of graphs, enabling tasks such as graph generation and reconstruction. This sort of method includes models like VAE [simonovsky2018graphvae] and GAN [wang2018graphgan].
  • Large Language Models: Recently, LLMs have been adapted for graph data by encoding graphs into tokens. This approach leverages the power of LLMs to process graphs, opening up new avenues for graph learning.
@@ 8.2 Graph Reinforcement Learning
Deep Reinforcement Learning (DRL) is a subfield of machine learning that integrates deep learning with reinforcement learning [landers2023deep]. DRL training agent to make decisions by learning optimal policy through interactions with the environment, using deep neural networks to handle high-dimensional state spaces and approximate complex value functions [milani2024explainable]. The sequential decision-making process can be modeled using a Markov Decision Process (MDP) or an extended MDP, such as a partially observable Markov decision process (POMDP) or Semi-MDP. Specifically, an MDP can be defined as $<S,A,T,R>$, where $S$ is state space, $A$ is the action space, $T$ is state transition probability, and $R$ is the reward value obtained after actor $a$ is performed. This approach enables the agent to maximize cumulative rewards over time in sequential decision-making tasks.
@@ 8.2.1 Graph Enhanced DRL
In multi-agent DRL, a group of agents cooperate or compete to achieve a common goal, where the key is to understand the mutual interplay between agents. This architecture is widely used in various tasks, such as traffic signal control [devailly2021ig], vehicle control systems [yang2023hierarchical], and network routing [xu2024graph]. The communication between agents provides information about the states and environments of other agents. Various attention-based methods have been proposed to address communication issues in multi-agent systems. In G2ANet [liu2020multi], the authors use a complete graph to model the relationships between agents in large-scale multi-agent scenarios and designed a two-stage attention model as the communication model. Similarly, in Graphocomm [shen2021graphcomm], the authors apply a relational graph module to model explicit and implicit relationships through explicit and implicit relation layers. Specifically, explicit relationships are built from prior knowledge, while implicit relationships are learned through interactions among agents. In SRI-AC [zhang2021structural], the authors deploy a Variational Autoencoder (VAE) model to predict interactions and learn a state representation from observational data.
@@ 8.4.1 Knowledge Extraction
Knowledge extraction (or knowledge acquisition) tasks can be divided into three categories: knowledge graph completion (KGC), relation extraction and entity discovery [DBLP:journals/tnn/JiPCMY22]. KGC aims to predict missing links in existing KGs, and embedding-based ranking, relation path reasoning, rule-based reasoning, and meta relational learning are main categories for KGC. Compared to KGC, relation extraction and entity discovery are for discovering new knowledge (e.g., relations and entities) from the text. Relation extraction models utilize attention mechanisms, GCNs, adversarial training, reinforcement learning, deep residual learning, and transfer learning. KGC, relation extraction, and entity discovery are three separate tasks for knowledge extraction. In order to simplify development and be convenient for knowledge extraction, proposing a unified framework is one of the challenges in this field. For example, DBLP:conf/aaai/Han0S18 propose a joint learning framework with mutual attention for data fusion between knowledge graphs and text, which simultaneously solves KGC and relation extraction from text. Additionally, knowledge extraction also face other challenges in KGC, relation extraction, and entity discovery as follows.

### v2

@@ 7.1.1 Differential Privacy Based Methods
xiang2024preserving presented a solution that focuses on node-level privacy through a two-component protocol: 1) Heter-Poisson, a specialized node sampling strategy that generates sub-graphs with desired properties, and 2) a randomization routine utilizing symmetric multivariate Laplace noise, offering a robust privacy guarantee. Additionally, the authors critiqued a prior approach [sajadmanesh2023gap] that utilizes differentially private node embeddings, highlighting fundamental privacy flaws and proving the impossibility of achieving both strong privacy and acceptable utility with such embeddings, thus underscoring intrinsic utility barriers in enforcing differential privacy.
@@ 7.1.3 Adversarial Privacy Preservation
7.2 Robustness of Graph Learning
Graph learning models are increasingly vulnerable to adversarial attacks, making the development of robust graph learning a critical area of research. These defense strategies are generally categorized into three main approaches: Adversarial Training, Graph Denoising, and Certifiable Robustness. The following sections introduce key methods within each category, along with other notable defense techniques that fall outside these classifications.
7.2.1 Adversarial Training
This approach involves fortifying a graph learning model by exposing it to adversarial samples during the training process, a practice inspired by its success in computer vision. The core idea is to improve the model’s ability to resist attacks by explicitly incorporating adversarial perturbations into the training loop. This can be achieved in two main ways:
Min-Max Optimization: A graph learning is trained to simultaneously minimize its loss on both clean and adversarially perturbed data. This process can be formally expressed by the following equation:
$\min_{\theta}\max_{\Delta_{\mathbf{A}}\in\mathcal{P}_{\mathbf{A}},\Delta_{\mathbf{x}}\in\mathcal{P}_{\mathbf{X}}}\mathcal{L}_{\text{train }}\left(f_{\theta}\left(\mathbf{A}+\Delta_{\mathbf{A}},\mathbf{X}+\Delta_{\mathbf{X}}\right)\right),$ (7.2)
where $\mathcal{L}_{\text{train}}$ is the classification loss on the labeled nodes. $\Delta_{\mathbf{A}}$ and $\Delta_{\mathbf{X}}$ represent the perturbations on the topology structure and node attributes, respectively. $\mathcal{P}_{\mathbf{A}}$ and $\mathcal{P}_{\mathbf{X}}$ denote the allowable perturbations within the attack budget.
This approach forces the model to learn a more robust decision boundary that is less susceptible to small, malicious changes in the input. For example, method like PerturbEmbedding adds an adversarial term to the loss function, encouraging the model to become more resistant to both structural and feature-based attacks. yang2025unifying proposed PerturbEmbedding that investigates the vulnerability of graph learning models to adversarial attacks that target both node features and graph structure. While acknowledging that existing adversarial training methods have shown promise in improving model robustness, they note that the application of these methods is often limited to specific datasets and model types. To address this limitation, they introduce a novel method, PerturbEmbedding, which applies adversarial perturbations directly to the hidden embeddings of graph learning models. This approach provides a unified and effective framework that enhances both the robustness and generalization of a wide range of graph learning models. By perturbing the hidden embeddings, the method forces the model to be resilient not just against input-level attacks, but also against malicious changes at intermediate layers, thereby creating a more robust and adaptable model.
Robust Pre-training: Some methods focus on pre-training the model on a large number of synthetically generated adversarial graphs before fine-tuning it on the clean graph. This allows the model to learn general robust features from a diverse set of attack scenarios. For example, a recent study proposes LLM4RGNN, which leverages the inference capabilities of LLMs to enhance a graph learning model’s robustness [zhang2025can]. This framework distills the knowledge of a powerful model like GPT-4 into a local LLM to identify and remove malicious edges from the graph structure. It also uses an additional language model to predict and recover important missing edges, effectively "denoising" the graph to make it more resilient to topology-based attacks.
7.2.2 Graph Denoising
The goal of graph denoising is to purify the input graph before or during the graph learning model’s forward pass, effectively neutralizing adversarial perturbations. This approach is based on the assumption that malicious perturbations are a form of noise that can be filtered out. Denoising methods typically focus on cleaning either the graph structure or the node features.
Structure Denoising: These methods aim to identify and remove suspicious or harmful links in the graph. The process can be formulated as a joint optimization problem to learn a clean adjacency matrix, $\mathbf{S}$, as shown below:
$\min_{\theta,\mathbf{S}}\mathcal{L}_{\text{train }}\left(f_{\theta}(\mathbf{X},\mathbf{S})\right)+\alpha\|\mathbf{A}-\mathbf{S}\|_{F}^{2}+\beta\|\mathbf{S}\|_{*}+\lambda\operatorname{tr}\left(\mathbf{X}^{T}\hat{\mathbf{L}}\mathbf{X}\right),$ (7.3)
where $\mathcal{L}_{\text{train}}(f_{\theta}(\mathbf{X},\mathbf{S}))$ is the classification loss. The term $|\mathbf{S}\|_{*}$ is the nuclear norm of the learned adjacency matrix $\mathbf{S}$, which encourages it to be low-rank. The final term, $\operatorname{tr}\left(\mathbf{X}^{T}\hat{\mathbf{L}}\mathbf{X}\right)$, encourages the learned adjacency matrix to form links between nodes with similar features, where $\hat{\mathbf{L}}$ is the Laplacian matrix of $\mathbf{S}$.
For instance, in real-world graphs, both structural noise (e.g., noisy edges) and a limited number of labeled nodes can significantly impair the performance of graph learning models. A recent framework proposed by dai2022towards addresses this by using noisy edges as a form of supervision to learn a denoised and dense graph. This process effectively eliminates or down-weights noisy edges while simultaneously creating new, more meaningful connections that facilitate the message-passing mechanism. The framework then uses these newly generated edges to regularize predictions for unlabeled nodes, improving the overall training and robustness of the graph learning model, even when faced with significant data imperfections.
Feature Denoising: This involves applying a denoising operation to the node features to remove adversarial noise. For instance, some approaches use a pre-processing step to project the node features onto a low-dimensional subspace, where the adversarial perturbations are assumed to be less significant. A novel framework, DiffSP (Diffusion-based Structure Purification), offers a prior-free approach to this problem [luo2025robust]. Instead of relying on assumptions about the attack, DiffSP uses a graph diffusion model to learn the inherent distribution of clean graphs. It then purifies a perturbed graph through a forward diffusion process (adding noise) and a reverse denoising process, where it removes adversarial noise under the guidance of the learned patterns. To avoid losing important information, DiffSP uses an LID-driven non-isotropic diffusion mechanism to selectively inject noise. The framework also promotes semantic alignment between the clean and purified graphs via a graph transfer entropy guided denoising mechanism, ultimately achieving robust graph learning without needing prior knowledge of the attack.
7.2.3 Certifiable Robustness
Unlike the first two categories, which provide empirical defense, certifiable robustness offers a mathematical guarantee that a graph learning model’s prediction will remain unchanged within a certain radius of perturbation. This provides a concrete measure of the model’s security. For a given node $v\in\mathcal{V}$ with label $y$, its robustness can be certified by solving the following optimization problem:
$m(v;\theta)=\min_{\hat{\mathcal{G}}\in\Phi(\mathcal{G})}\max_{i\neq y}f_{\theta}(\hat{\mathcal{G}})_{vy}-f_{\theta}(\hat{\mathcal{G}})_{vi},$ (7.4)
where $f_{\theta}(\hat{\mathcal{G}})_{vi}$ is the predicted logit of node $v$ in class $i$, and $\Phi(\mathcal{G})$ includes all allowable perturbed versions of the graph. A model is certifiably robust for node $v$ if $m(v)>0$. Key methods in this area include:
Randomized Smoothing: This technique, adapted from computer vision, involves adding random noise to the input graph and averaging the model’s predictions over many noisy samples. By doing so, it provides a provable robustness guarantee for the model’s output, ensuring that the prediction remains consistent as long as the perturbation is within a specified bound. Building on this idea, the node-aware bi-smoothing framework extends this technique to address a new challenge: certified robustness against graph injection attacks (GIAs) [lai2024node]. This framework is the first to provide a verifiable defense for general node classification tasks against GIAs, making it applicable for both evasion and poisoning attacks. The method is also model-agnostic, meaning it can be used with various graph learning architectures. The authors of this work provide a rigorous theoretical analysis to establish the certifiable conditions of their smoothing scheme and demonstrate its practical effectiveness as a defense against real-world attacks.
Interval Bound Propagation (IBP): This approach, inspired by IBP methods for neural networks, aims to compute a tight bound on a graph learning model’s output for all possible inputs within a given perturbation set. By ensuring that this output range does not cross a decision boundary, it can certify the model’s robustness against a wide range of attacks. A recent study addresses this challenge by applying a dual backpropagation network to deep graph learning [zeng2025robustness]. The authors propose a novel method for calculating intermediate activation bounds based on a linear approximation, which effectively enhances verification accuracy and scalability. Their approach is shown to be capable of verifying the robustness of deep graph learning with up to $32$ layers, demonstrating a significant improvement in the maximum perturbation value that correctly classified nodes can withstand.
In addition to these three main categories, other notable defense methods are emerging, such as those that leverage graph diffusion to disperse adversarial information or knowledge distillation to transfer robustness from a large, robust model to a smaller one. The continuous evolution of attack and defense strategies highlights the dynamic and critical nature of building trustworthy and robust graph learning systems.
@@ 7.3.1 Group Fairness
Graph learning can produce biased predictions based on protected sensitive attributes, such as skin color and gender, due to historical biases in the training data. This discrimination can be exacerbated by the graph structures and the message-passing mechanisms inherent in graph learning, particularly in high-stakes applications like crime rate prediction. While extensive research has focused on fair classification with independently and identically distributed (i.i.d.) data, approaches for non-i.i.d. scenarios remain limited, especially since sensitive attributes are often scarce in real-world graphs. To tackle this issue, FairGNN [dai2023learning] has been proposed to mitigate bias while maintaining node classification accuracy by leveraging limited sensitive information, and it employs local differential privacy to protect sensitive attributes. An extension, NT-FairGNN, further addresses the challenges of achieving fairness and privacy simultaneously, with theoretical analysis and empirical results demonstrating their effectiveness.
@@ 8.1 Graph Foundation Models
Graph Foundation Models (GFMs) are large-scale pre-trained models specifically designed for processing graph-structured data. These models are generally pre-trained on vast amounts of data across various domains [xia2022asurvey, lengthiu2023towardsgraph]. Besides, GFMs can adapt to varying demands with minimal fine-tuning or even in few-shot/zero-shot scenarios. There exist many open issues to be addressed, and we discuss them from three distinct perspectives: pre-training, architecture design, and adaptation strategies.
@@ 8.1.2 Architecture Design
  • Message-Passing Models: Message-passing models are the most common designs in graph models. These architectures capture local graph structures by iteratively passing and aggregating information between nodes. Classic models in this category include GCN [KipfW17gcn], GAT [velickovic2018graph], and GraphSAGE [hamiltonYL17graphsage]. Despite the success, message-passing struggles with expressivity, often failing to capture long-range dependencies due to over-smoothing and limited receptive fields.
  • Graph Transformer: Inspired by the success of transformers in natural language processing, Graph Transformers (GTs) [ying2021transformers, chen2022structure, ma2024polyformer] have been extended to graph-structured data. These models apply global attention mechanisms to capture long-range dependencies in graphs, making them well-suited for handling complex relations. Their token-based graph representations are well-suited for pretraining tasks, although scalability and structural priors remain open challenges.
  • Generative Models: Generative graph models, such as VAEs [simonovsky2018graphvae] and GAN-based methods [wang2018graphgan], learn latent graph distributions to support tasks like generation, completion, and anomaly detection. Besides, recent diffusion and autoregressive approaches [kong2023autoregressive] enable high-quality graph synthesis, capturing structural and semantic priors that enhance the versatility of Graph Foundation Models through generative pretraining.
  • Large Language Models: Recently, LLMs have been adapted for graph data by encoding graphs into tokens [mao2024position, tang2024graphgpt]. This approach leverages the power of LLMs to process graphs, opening up new avenues for graph learning. While promising, this line of work raises open questions on how to design good tokenization schemes, inject structural priors, and mitigate hallucination. As GFMs evolve, the convergence of LLM-based approaches and GTs may play a central role in more areas.
@@ 8.1.3 Tuning, Adaptation and Compression
GFMs are emerging as a unified framework for graph-based learning, combining large-scale pretraining, GTs, and flexible fine-tuning to support diverse applications from social networks to scientific discovery. As the field matures, GFMs may become core components of graph learning and become a unified tool across graph tasks.
@@ 8.2 Graph Reinforcement Learning
Deep Reinforcement Learning (DRL) is a subfield of machine learning that integrates deep learning with reinforcement learning [landers2023deep]. DRL training agent to make decisions by learning optimal policy through interactions with the environment, using deep neural networks to handle high-dimensional state spaces and approximate complex value functions [milani2024explainable]. The sequential decision-making process can be modeled using a Markov Decision Process (MDP) or an extended MDP, such as a partially observable Markov decision process (POMDP) or Semi-MDP. Specifically, an MDP can be defined as <S,A,T,R><S,A,T,R>, where $S$ is state space, $A$ is the action space, $T$ is state transition probability, and $R$ is the reward value obtained after actor $a$ is performed. This approach enables the agent to maximize cumulative rewards over time in sequential decision-making tasks.
@@ 8.2.1 Graph Enhanced DRL
In multi-agent DRL, a group of agents cooperate or compete to achieve a common goal, where the key is to understand the mutual interplay between agents. This architecture is widely used in various tasks, such as traffic signal control [devailly2021ig], vehicle control systems [yang2023hierarchical], and network routing [xu2024graph]. The communication between agents provides information about the states and environments of other agents. Various attention-based methods have been proposed to address communication issues in multi-agent systems. In G2ANet [liu2020multi], the authors used a complete graph to model the relationships between agents in large-scale multi-agent scenarios and designed a two-stage attention model as the communication model. Similarly, in Graphocomm [shen2021graphcomm], the authors applied a relational graph module to model explicit and implicit relationships through explicit and implicit relation layers. Specifically, explicit relationships are built from prior knowledge, while implicit relationships are learned through interactions among agents. In SRI-AC [zhang2021structural], the authors deployed a Variational Autoencoder (VAE) model to predict interactions and learn a state representation from observational data.
@@ 8.4.1 Knowledge Extraction
Knowledge extraction (or knowledge acquisition) tasks can be divided into three categories: knowledge graph completion (KGC), relation extraction, and entity discovery [DBLP:journals/tnn/JiPCMY22]. KGC aims to predict missing links in existing KGs, and embedding-based ranking, relation path reasoning, rule-based reasoning, and meta relational learning are main categories for KGC. Compared to KGC, relation extraction and entity discovery are for discovering new knowledge (e.g., relations and entities) from the text. Relation extraction models utilize attention mechanisms, GCNs, adversarial training, reinforcement learning, deep residual learning, and transfer learning. KGC, relation extraction, and entity discovery are three separate tasks for knowledge extraction. In order to simplify development and be convenient for knowledge extraction, proposing a unified framework is one of the challenges in this field. For example, DBLP:conf/aaai/Han0S18 proposed a joint learning framework with mutual attention for data fusion between knowledge graphs and text, which simultaneously solves KGC and relation extraction from text. Additionally, knowledge extraction also face other challenges in KGC, relation extraction, and entity discovery as follows.